# xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology

### v1

@@ Abstract
Multiple instance learning (MIL) is an effective and widely used approach for weakly supervised machine learning. In histopathology, MIL models have achieved remarkable success in tasks like tumor detection, biomarker prediction, and outcome prognostication. However, MIL explanation methods are still lagging behind, as they are limited to small bag sizes or disregard instance interactions. We revisit MIL through the lens of explainable AI (XAI) and introduce xMIL, a refined framework with more general assumptions. We demonstrate how to obtain improved MIL explanations using layer-wise relevance propagation (LRP) and conduct extensive evaluation experiments on three toy settings and four real-world histopathology datasets. Our approach consistently outperforms previous explanation attempts with particularly improved faithfulness scores on challenging biomarker prediction tasks. Finally, we showcase how xMIL explanations enable pathologists to extract insights from MIL models, representing a significant advance for knowledge discovery and model debugging in digital histopathology. [^1] We are pleased to share our codes upon request.
@@ 1 Introduction
Explaining which visual features a MIL model uses for its prediction is highly relevant in this context. It allows experts to sanity-check the model strategy [19], e.g., whether a model focuses on the disease area for making a diagnosis. This is particularly important in histopathology, where models operating in high-stake environments are prone to learning confounding factors like artifacts or staining differences instead of actual signal [20, 21, 22]. On top of that, MIL explanations can enable pathologists to discover novel connections between visual features and prediction targets. For example, the explanations could reveal a previously unknown association of a histopathological pattern with poor survival, leading to the identification of a targetable disease mechanism. Previous works have shown the potential of scientific knowledge discovery from ML explanation methods [22, 23, 24, 25, 26].
@@
We revisit MIL through the lens of explainable AI (XAI) and introduce xMIL, a more general and realistic multiple instance learning framework including requirements for good explanations. We then present xMIL-LRP, an adaptation of layer-wise relevance propagation (LRP [35, 36]) to MIL. xMIL-LRP distinguishes between positive and negative evidence, disentangles instance interactions, and scales to large bag sizes. It applies to various MIL models without requiring architecture modifications, including Attention MIL [3] and TransMIL [4]. We assess the performance of multiple explanation techniques via three toy experiments, which can serve as a novel benchmarking tool for MIL explanations in complex tasks with instance interactions and context-sensitive targets. We further perform faithfulness experiments on four real-world histopathology datasets covering tumor detection, disease subtyping, and biomarker prediction. xMIL-LRP consistently outperforms previous attempts across all tasks and model architectures, with the biggest advantages observed for Transformer-based biomarker prediction.
@@
The paper is structured as follows: In Section 2, we review MIL assumptions, models, and explanation methods related to this work. In Section 3, we introduce xMIL as a general form of MIL, and xMIL-LRP as a solution for it. In Section 4, we experimentally show the improved explanation quality of our approach. We demonstrate how to extract insights from example heatmaps in Section 5.
@@ 2.1 Multiple instance learning (MIL)
MIL models. MIL architectures typically consist of three components as illustrated in Fig. 2: a backbone extracting instance representations, an aggregation function fusing the instance representations into a bag representation, and a prediction head inferring the final bag prediction. As recent foundation models for histopathology have become powerful feature extractors suitable for a wide range of tasks [29, 39, 40, 41, 42], the weights of the backbone are often frozen, allowing for a more efficient training. For aggregation, earlier works used parameter-free mean or max pooling approaches [43, 44, 45]. Recently, attention mechanisms could improve performance, flexibly extracting relevant instance-level information using non-linear weighting [3, 6, 46] and self-attention [4, 47]. Attention MIL (AttnMIL) [3] computes a weighted average of the instances’ feature vectors via a single attention head. TransMIL [4] uses a custom two-layer Transformer architecture, viewing instance representations as tokens. The bag representation is extracted from the class token at the final layer. TransMIL allows for computing arbitrary pairwise interactions between all instances relevant to the prediction task. While various extensions of AttnMIL and TransMIL have been proposed (e.g., [5, 6, 7, 8, 9, 10, 11, 12]), these two methods are arguably prototypical and among the most commonly used in the digital histopathology community.
@@ 3 Methods
Figure 2: The two steps of xMIL: estimating the aggregation function (A) and the evidence function (B). Panel A shows a block diagram of a MIL model applied to a histopathology slide. The feature extraction module is typically a combination of a frozen foundation model followed by a shallow MLP. In most of the recent MIL models, the aggregation module uses attention mechanisms for combining the instance feature vectors into a single feature representation per bag. The prediction head is a linear layer or an MLP. Panel B schematically shows xMIL-LRP for explaining AttnMIL. In xMIL-LRP, the model output is backpropagated to the input instances. The colored lines represent the relevance flow, with red indicating positive values and blue negative ones. The attention module is handled via the AH-rule as described in Section 3.2. As discussed in Section 3.3, the instance explanation scores can be computed at the output of the foundation model or at the input level.
@@
We introduce xMIL-LRP as an efficient solution to xMIL, bringing layer-wise relevance propagation (LRP) to MIL. LRP is a well-established XAI method [35, 52] with a large body of literature supporting its performance in explaining various types of architectures in different tasks [32, 36, 53, 54, 55, 56]. Starting from the prediction score of a selected class, the LRP attribution of neuron $i$ in layer $l$ receives incoming messages from neurons $j$ from subsequent layer $l+1$, resulting in relevance scores $R_{i}^{(l)}=\sum_{j}\frac{q_{ij}}{\sum_{i^{\prime}}q_{i^{\prime}j}}\cdot R_{j}^{(l+1)}$, with $q_{ij}$ being the contribution of neuron $i$ of layer $l$ relevance $R_{j}^{(l+1)}$. A variety of so-called “propagation rules” have been proposed [36] to specify the contribution $q_{ij}$ in specific model layers. For the attention mechanism, as a core component of many MIL architectures, we employ the AH-rule introduced by Ali et al. [32], which allocates feature-wise relevance scores to tokens from the preceding layer based on attention weights as $R_{i}^{(l)}=\sum_{j}\frac{x_{i}p_{ij}}{\sum_{i^{\prime}}x_{i^{\prime}}p_{i^{\prime}j}}\cdot R_{j}^{(l+1)}$, where $x_{i}$ is the activation of a neuron $i$, and $p_{ij}$ an attention score between neurons $j$ and $i$. We illustrate the effect of this rule in AttnMIL in Figure 2-B. The relevance flow separates the instances weighted by the attention mechanism into positive, negative, and neutral instances, resulting in more descriptive heatmaps that better show the relevant tissue regions compared to attention scores.
We further implement the LRP-$\gamma$ rule for linear layers followed by ReLU activation function [52], as well as the LN-rule to address the break of conservation in layer norm [32]. At the instance-level, xMIL-LRP assigns each instance $\mathbf{x}_{k}$ a relevance vector $\mathbf{r}_{k}=[R_{k,1}^{(0)},\cdots,R_{k,D}^{(0)}]^{T}$ with $R_{k,d}^{(0)}$ being the relevance score of the $d$-th feature of $\mathbf{x}_{k}\in\mathbb{R}^{D}$. We define the instance-wise relevance score as an estimate for the evidence score of the instance as $\hat{\epsilon}_{k}=\sum_{d}R_{k,d}^{(0)}$.
@@ 3.3 Properties of xMIL-LRP and other explanation methods
Context sensitivity: xMIL-LRP disentangles instance interactions and contextual information as it jointly considers the relevance flow across the whole bag. LRP and GxI are rooted in a deep Taylor decomposition of the model prediction [57] and consequently capture dependencies between features by tracing relevance flow through the components of the MIL model. While attention is context-aware, it is limited to considering dependencies of features at a specific layer. The “single” method is unaware of context. “One-removed” and additive MIL can only capture the impact of individual instances on the prediction.
@@
Conservation: Following the conservation principle of LRP, xMIL-LRP provides an instance-wise decomposition of the model output, i.e., $\sum_{k}\hat{\epsilon}_{k}=\sum_{k,d}R^{(0)}_{k,d}=y$. This instance-level conservation also holds for additive MIL, but not for the other discussed methods. The local conservation principle of LRP [36] enables the analysis of attribution scores at the level of instance feature vectors and further provides a starting point for propagation through the foundation model—the sum of attribution scores is consistent across layers of the model.
Pixel-level explanations: By propagating the feature-wise relevance scores at the output of the foundation model through the vision backbone of our MIL pipeline, we can obtain input-level relevance scores. Therefore, the estimated instance evidence scores at the patch level can be further explained by computing pixel-level heatmaps, allowing a detailed analysis of relevant visual features [49]. While computing pixel-level explanations is computationally infeasible for perturbation-based approaches, they are generally not defined in attention-based explanation methods and additive models.
@@ 4.1 Toy experiments
We designed novel toy experiments to assess and compare the characteristics of xMIL-LRP and the baseline methods in controlled settings. We focused on evaluating to what extent the explanations account for context sensitivity and positive and negative evidence, i.e., the first two characteristics of the evidence function according to Definition 3.2, which we consider crucial aspects for explaining real-world histopathology prediction tasks.
@@
Table 1: Results of the toy experiments. We report AUPRC-2 scores of MIL explanation methods on three toy datasets measuring how well a method identified instances with positive and negative evidence scores (mean ± std. over 10 repetitions). The highest mean scores are bold and the second highest ones are underlined. We also display the model performances ("Test AUROC", mean ± std.)
@@
AttnMIL TransMIL AttnMIL TransMIL AttnMIL TransMIL
Test AUROC 0.995 ± 0.001 0.997 ± 0.001 0.971 ± 0.001 0.981 ± 0.003 0.884 ± 0.082 0.918 ± 0.058
Rand 0.314 ± 0.001 0.311 ± 0.003 0.419 ± 0.002 0.416 ± 0.003 0.539 ± 0.003 0.536 ± 0.002
Attn 0.534 ± 0.001 0.522 ± 0.004 0.448 ± 0.003 0.455 ± 0.030 0.614 ± 0.006 0.602 ± 0.012
Single 0.869 ± 0.017 0.845 ± 0.071 0.884 ± 0.004 0.913 ± 0.021 0.732 ± 0.063 0.764 ± 0.085
GxI 0.717 ± 0.080 0.404 ± 0.064 0.715 ± 0.144 0.443 ± 0.055 0.634 ± 0.045 0.574 ± 0.054
xMIL-LRP 0.909 ± 0.011 0.904 ± 0.021 0.906 ± 0.008 0.920 ± 0.011 0.768 ± 0.037 0.811 ± 0.043
To assess the explanation quality, we first defined valid evidence scores as ground truths according to Definition 3.2. For each dataset, we require one evidence function per predicted class $c$, denoted by $\mathcal{E}^{(c)}(X,y,\mathbf{x}_{k})=\epsilon_{k}^{(c)}$. We assigned $\epsilon_{k}^{(c)}=1$ if $\textbf{x}_{k}$ supports class $c$, $\epsilon_{k}^{(c)}=-1$ if the instance refutes class $c$, and $\epsilon_{k}^{(c)}=0$ if it is irrelevant. We aimed to measure whether an explanation method correctly distinguishes instances with positive, neutral, and negative evidence scores. Therefore, we computed a two-class averaged area under the precision-recall curve (AUPRC-2), measuring if the positive instances received the highest and the negative instances the lowest explanation scores. We assessed AttnMIL and TransMIL models and repeated each experiment 10 times. The details of the ground truth, the evaluation metric, and the experimental setup are provided in Appendix A.2.
Table 1 displays the test AUROC scores of the two models across datasets, demonstrating that the models solve the tasks successfully, though to varying degrees, along with the performances of the explanation methods. We find that xMIL-LRP consistently outperformed other explanation approaches across MIL models and datasets. It reached particularly high AUPRC-2 scores in the 4-Bags and Pos-Neg datasets while being most robust in the more difficult Adjacent Pairs setting. Attention suffered significantly from the presence of positive and negative evidence, which it cannot distinguish by design. The single perturbation method provided good explanations in the Pos-Neg setting, where numbers have a fixed evidence score irrespective of the other instances in the bag. However, in 4-Bags and Adjacent Pairs the method’s performance decreased, as it always assigns the same score to the same instance regardless of the bag context. In contrast, xMIL-LRP is both context-sensitive and identifies positive and negative instances. Since these aspects are common features of many real-world histopathological datasets, we conclude that our method is the only suitable approach for such complex settings.

### v2

@@ Abstract
Multiple instance learning (MIL) is an effective and widely used approach for weakly supervised machine learning. In histopathology, MIL models have achieved remarkable success in tasks like tumor detection, biomarker prediction, and outcome prognostication. However, MIL explanation methods are still lagging behind, as they are limited to small bag sizes or disregard instance interactions. We revisit MIL through the lens of explainable AI (XAI) and introduce xMIL, a refined framework with more general assumptions. We demonstrate how to obtain improved MIL explanations using layer-wise relevance propagation (LRP) and conduct extensive evaluation experiments on three toy settings and four real-world histopathology datasets. Our approach consistently outperforms previous explanation attempts with particularly improved faithfulness scores on challenging biomarker prediction tasks. Finally, we showcase how xMIL explanations enable pathologists to extract insights from MIL models, representing a significant advance for knowledge discovery and model debugging in digital histopathology. Codes are available at: https://github.com/tubml-pathology/xMIL.
@@ 1 Introduction
Explaining which visual features a MIL model uses for its prediction is highly relevant in this context. It allows experts to sanity-check the model strategy [19], e.g., whether a model focuses on the disease area for making a diagnosis. This is particularly important in histopathology, where models operating in high-stake environments are prone to learning confounding factors like artifacts or staining differences instead of actual signal [20, 21, 22]. On top of that, MIL explanations can enable pathologists to discover novel connections between visual features and prediction targets. For example, the explanations could reveal a previously unknown association of a histopathological pattern with poor survival, leading to the identification of a targetable disease mechanism. Previous works have shown the potential of scientific knowledge discovery from explainable AI (XAI) [22, 23, 24, 25, 26].
@@
We revisit MIL through the lens of XAI and introduce xMIL, a more general and realistic multiple instance learning framework including requirements for good explanations. We then present xMIL-LRP, an adaptation of layer-wise relevance propagation (LRP) [35, 36] to MIL. xMIL-LRP distinguishes between positive and negative evidence, disentangles instance interactions, and scales to large bag sizes. It applies to various MIL models without requiring architecture modifications, including Attention MIL [3] and TransMIL [4]. We assess the performance of multiple explanation techniques via three toy experiments, which can serve as a novel benchmarking tool for MIL explanations in complex tasks with instance interactions and context-sensitive targets. We further perform faithfulness experiments on four real-world histopathology datasets covering tumor detection, disease subtyping, and biomarker prediction. xMIL-LRP consistently outperforms previous attempts across all tasks and model architectures, with the biggest advantages observed for Transformer-based biomarker prediction.
@@
The paper is structured as follows: In Section 2, we review MIL assumptions, models, and explanation methods related to this work. In Section 3, we introduce xMIL as a general form of MIL, and xMIL-LRP as a solution for it. In Section 4, we experimentally show the improved explanation quality of our approach. We demonstrate how to extract insights from example heatmaps in Section 5. Our contributions are summarized as follows:
  • Methodical: Despite attempts to apply XAI to MIL models in histopathology (e.g. [6, 27, 28, 29, 33, 34, 37, 38, 39, 40]), there exists no formalism guiding the interpretation of the heatmaps and defining their desired properties. xMIL is a novel framework addressing this gap. Within xMIL, heatmaps estimate the instances’ impact on the bag label, which makes their interpretation straightforward and insightful.
  • Empirical: Our extensive empirical evaluation of XAI methods for MIL on synthetic and real-world histopathology datasets is the first of its kind. It reveals that the widely used MIL explanation methods regularly yield misleading results. In contrast, xMIL-LRP sets a new state-of-the-art for explainability in AttnMIL and TransMIL models in histopathology.
  • Insight generation: Previous studies [33, 34] conducted qualitative assessments of heatmaps on easy-to-learn datasets like CAMELYON or TCGA NSCLC. The insights gained in these settings are limited to model debugging, i.e., “Does the model focus on the disease area?” To our knowledge, we are the first to present a method generating heatmaps that enable pathologists to extract fine-grained insights about the model in a difficult biomarker prediction task.
@@ 2.1 Multiple instance learning (MIL)
MIL models. MIL architectures typically consist of three components as illustrated in Figure 2: a backbone extracting instance representations, an aggregation function fusing the instance representations into a bag representation, and a prediction head inferring the final bag prediction. As recent foundation models for histopathology have become powerful feature extractors suitable for a wide range of tasks [29, 43, 44, 45, 46], the weights of the backbone are often frozen, allowing for a more efficient training. For aggregation, earlier works used parameter-free mean or max pooling approaches [47, 48, 49]. Recently, attention mechanisms could improve performance, flexibly extracting relevant instance-level information using non-linear weighting [3, 6, 50] and self-attention [4, 51]. Attention MIL (AttnMIL) [3] computes a weighted average of the instances’ feature vectors via a single attention head. TransMIL [4] uses a custom two-layer Transformer architecture, viewing instance representations as tokens. The bag representation is extracted from the class token at the final layer. TransMIL allows for computing arbitrary pairwise interactions between all instances relevant to the prediction task. While various extensions of AttnMIL and TransMIL have been proposed (e.g., [5, 6, 7, 8, 9, 52, 53, 10, 11, 12]), these two methods are arguably prototypical and among the most commonly used in the digital histopathology community.
@@ 3 Methods
Figure 2: The two steps of xMIL: estimating the aggregation function (A) and the evidence function (B). Panel A shows a block diagram of a MIL model applied to a histopathology slide. The feature extraction module is typically a combination of a frozen foundation model followed by a shallow MLP. In most of the recent MIL models, the aggregation module uses attention mechanisms for combining the instance feature vectors into a single feature representation per bag. The prediction head is a linear layer or an MLP. Panel B schematically shows xMIL-LRP for explaining AttnMIL. In xMIL-LRP, the model output is backpropagated to the input instances. The colored lines represent the relevance flow. Red and blue colors encode the positive and negative values. The attention module is handled via the AH-rule as described in Section 3.2. As discussed in Section 3.3, the instance explanation scores can be computed at the output of the foundation model or at the input level.
Notation. We denote vectors with boldface lowercase letters (e.g., $\mathbf{x}$), scalars with lowercase letters (e.g., $x$), and sets with uppercase letters (e.g., $X$).
@@
We introduce xMIL-LRP as an efficient solution to xMIL, bringing layer-wise relevance propagation (LRP) to MIL. LRP is a well-established XAI method [35, 58] with a large body of literature supporting its performance in explaining various types of architectures in different tasks [32, 36, 59, 60, 61, 62]. Starting from the prediction score of a selected class, the LRP attribution of neuron $i$ in layer $l$ receives incoming messages from neurons $j$ from subsequent layer $l+1$, resulting in relevance scores $r_{i}^{(l)}=\sum_{j}\frac{q_{ij}}{\sum_{i^{\prime}}q_{i^{\prime}j}}\cdot r_{j}^{(l+1)}$, with $q_{ij}$ being the contribution of neuron $i$ of layer $l$ to relevance $r_{j}^{(l+1)}$  [^1] In some XAI research papers, uppercase letters, e.g., $R_{j}^{(l)}$, are used for denoting relevance values.. A variety of so-called “propagation rules” have been proposed [36] to specify the contribution $q_{ij}$ in specific model layers. For the attention mechanism, as a core component of many MIL architectures, we employ the AH-rule introduced by Ali et al. [32]. In a general attention mechanism, let $\mathbf{z}_{k}=\left[z_{kd}\right]_{d}$ be the embedding vector of the $k$-th token and $p_{kj}$ the attention score between tokens $k$ and $j$. The output vector of the attention module is $\mathbf{y}_{j}=\sum_{k}p_{kj}\mathbf{z}_{k}$. The AH-rule of LRP treats attention scores as a constant weighting matrix during the backpropagation pass of LRP. If $R(y_{jd})$ is the relevance of the $d$-th dimension of $\mathbf{y}_{j}=\left[y_{jd}\right]_{d}$, the AH-rule computes the relevance of the $d$-th feature of $\mathbf{z}_{k}$ as:
$R(z_{kd})=\sum_{j}\frac{z_{kd}p_{kj}}{\sum_{i}z_{id}p_{ij}}R(y_{jd}).$ (1)
This formulation can be directly applied to AttnMIL, and also adapted to a QKV attention block in a transformer, where $\mathbf{z}_{k}$ is the embedding associated with the value representation.
We illustrate the effect of this rule in AttnMIL in Figure 2-B. The relevance flow separates the instances weighted by the attention mechanism into positive, negative, and neutral instances, resulting in more descriptive heatmaps that better show the relevant tissue regions compared to attention scores.
We further implement the LRP-$\epsilon$ rule for linear layers followed by ReLU activation function [58], as well as the LN-rule to address the break of conservation in layer norm [32], with details presented in Appendix A.2.
At the instance-level, xMIL-LRP assigns each instance $\mathbf{x}_{k}=[x_{kd}]_{d}\in\mathbb{R}^{D}$ a relevance vector $\mathbf{r}_{k}=[r_{kd}]_{d}$ with $r_{kd}=R(x_{kd})=r^{(0)}_{kd}$ being the relevance score of the $d$-th feature of $\mathbf{x}_{k}$. We define the instance-wise relevance score as an estimate for the evidence score of the instance as $\hat{\epsilon}_{k}=\sum_{d}r_{kd}$.
@@ 3.3 Properties of xMIL-LRP and other explanation methods
Context sensitivity: xMIL-LRP disentangles instance interactions and contextual information as it jointly considers the relevance flow across the whole bag. LRP and Gradient $\times$ Input (G×$\times$×I) are rooted in a deep Taylor decomposition of the model prediction [63] and consequently capture dependencies between features by tracing relevance flow through the components of the MIL model. While attention is context-aware, it is limited to considering dependencies of features at a specific layer. The “single” method is unaware of context. “One-removed” and additive MIL can only capture the impact of individual instances on the prediction.
@@
Conservation: Following the conservation principle of LRP, xMIL-LRP provides an instance-wise decomposition of the model output, i.e., $\sum_{k}\hat{\epsilon}_{k}=\sum_{k,d}r_{kd}=y$. This instance-level conservation also holds for additive MIL, but not for the other discussed methods. The local conservation principle of LRP [36] further allows us to analyze attribution scores at the instance feature vector level without requiring propagation through the foundation model—the instance-wise attribution scores are the same at any layer of the model.
@@ 4.1 Toy experiments
We designed novel toy experiments to assess and compare the characteristics of xMIL-LRP and the baseline methods for AttnMIL, TransMIL, and AddMIL in controlled settings. We focused on evaluating to what extent the explanations account for context sensitivity and positive and negative evidence, i.e., the first two characteristics of the evidence function according to Definition 3.2, which we consider crucial aspects for explaining real-world histopathology prediction tasks.
@@
To assess the explanation quality, we first defined valid evidence scores as ground truths according to Definition 3.2. For each dataset, we require one evidence function per predicted class $c$, denoted by $\mathcal{E}^{(c)}(X,y,\mathbf{x}_{k})=\epsilon_{k}^{(c)}$. We assigned $\epsilon_{k}^{(c)}=1$ if $\mathbf{x}_{k}$ supports class $c$, $\epsilon_{k}^{(c)}=-1$ if the instance refutes class $c$, and $\epsilon_{k}^{(c)}=0$ if it is irrelevant. We aimed to measure whether an explanation method correctly distinguishes instances with positive, neutral, and negative evidence scores. Therefore, we computed a two-class averaged area under the precision-recall curve (AUPRC-2), measuring if the positive instances received the highest and the negative instances the lowest explanation scores. We assessed AttnMIL and TransMIL models and repeated each experiment 30 times. The details of the ground truth, the evaluation metric, and the experimental setup are provided in Appendix A.3.
Table 1 displays the test AUROC scores of the three models across datasets, demonstrating that the models solve the tasks to varying degrees, alongside the performances of the explanation methods. We find that xMIL-LRP outperformed the other explanation approaches across MIL models and datasets in all but one setting. It reached particularly high AUPRC-2 scores in the 4-Bags and Pos-Neg datasets while being most robust in the more difficult Adjacent Pairs setting. Attention severely suffered from the presence of positive and negative evidence, which it cannot distinguish by design. While IG performed comparably to xMIL-LRP for AttnMIL models, it was inferior for TransMIL. Notably, the test AUROC of AddMIL was worse in all settings, resulting in explanations that are not competitive with the post-hoc explanation methods on AttnMIL and TransMIL. This supports our point that AddMIL may not perform competitively in difficult prediction tasks. The single perturbation method provided good explanations in the Pos-Neg setting, where numbers have a fixed evidence score irrespective of the other instances in the bag. However, in 4-Bags and Adjacent Pairs, the method’s performance decreased, as it always assigns the same score to the same instance regardless of the bag context. In contrast, xMIL-LRP is both context-sensitive and identifies positive and negative instances. Since we expect that these aspects are common features of many real-world histopathological datasets, we conclude that our method is the only suitable approach for such complex settings.
Table 1: Results of the toy experiments. We report AUPRC-2 scores of MIL explanation methods on three toy datasets measuring how well a method identified instances with positive and negative evidence scores (mean ± std. over 30 repetitions). The highest mean scores are bold and the second highest are underlined. We also display the model performances ("Test AUROC", mean ± std.).
@@
AttnMIL TransMIL AddMIL AttnMIL TransMIL AddMIL AttnMIL TransMIL AddMIL
Test AUROC 1.00 ± 0.00 1.00 ± 0.00 0.98 ± 0.02 0.97 ± 0.00 0.98 ± 0.00 0.89 ± 0.03 0.88 ± 0.08 0.92 ± 0.06 0.77 ± 0.07
Rand 0.31 ± 0.00 0.31 ± 0.00 0.42 ± 0.00 0.42 ± 0.00 0.54 ± 0.00 0.54 ± 0.00
Attn 0.53 ± 0.00 0.52 ± 0.00 0.54 ± 0.01 0.45 ± 0.00 0.46 ± 0.03 0.48 ± 0.02 0.61 ± 0.01 0.60 ± 0.01 0.63 ± 0.04
Single 0.87 ± 0.02 0.85 ± 0.07 0.89 ± 0.00 0.91 ± 0.02 0.73 ± 0.06 0.77 ± 0.06
Logits 0.79 ± 0.11 0.68 ± 0.17 0.71 ± 0.09
G×$\times$×I 0.72 ± 0.08 0.40 ± 0.07 0.72 ± 0.14 0.44 ± 0.06 0.63 ± 0.05 0.57 ± 0.05
IG 0.88 ± 0.01 0.80 ± 0.09 0.93 ± 0.00 0.82 ± 0.08 0.75 ± 0.03 0.72 ± 0.07
xMIL-LRP 0.91 ± 0.01 0.90 ± 0.02 0.91 ± 0.01 0.92 ± 0.01 0.77 ± 0.04 0.81 ± 0.04